The manager of the lodge has found that about 10% of the parties who make reservations
do not show up. If 120 party observations have been made,
(i) How many can be expected to show up?
(ii) Find the variance of this distribution.
(iii) Find the standard deviation of this distribution.
1
Expert's answer
2020-05-04T19:53:32-0400

Probability that party show up: P=100-10=90%=0.9

1) E(x)=n*p=120*0.9=108

2)Var(x)=npq=120*0.9*0.1=10.8

3)σ= "\\sqrt{{Var(x)}}= \u200b" 3.2863